### v3

# Neural Ordinary Differential Equations for Mapping the Magnetic QCD Phase Diagram via Holography

Abstract
The QCD phase diagram is crucial for understanding strongly interacting matter under extreme conditions, with major implications for cosmology, neutron stars, and heavy-ion collisions. We present a novel holographic QCD model utilizing neural ordinary differential equations (ODEs) to map the QCD phase diagram under magnetic field $B$, baryon chemical potential $\mu_{B}$, and temperature $T$. By solving the inverse problem of constructing the gravitational theory from Lattice QCD data, we reveal an unprecedentedly rich phase structure at finite $B$, including discovering multiple critical endpoints (CEPs) under strong magnetic fields. Specifically, for $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, we identify two distinct CEPs at $(T_{C}=87.3\,\mathrm{MeV},\,\mu_{C}=115.9\,\mathrm{MeV})$ and $(T_{C}=78.9\,\mathrm{MeV},\,\mu_{C}=244.0\,\mathrm{MeV})$. Notably, the critical exponents vary depending on the CEP’s location. These findings significantly advance our understanding of the QCD phase structure and provide concrete predictions for experimental validation at upcoming facilities such as FAIR, JPARC-HI, and NICA.
Introduction—Quantum Chromodynamics (QCD), the theory of strong interactions, governs quark and gluon behavior under extreme conditions, underpinning high-energy physics, astrophysics, and cosmology. The QCD phase diagram, which maps transitions between the deconfined quark-gluon plasma and hadronic matter, is central to understanding phenomena in non-central heavy-ion collisions, neutron stars, and the early universe Braun-Munzinger:2008szb ; Philipsen:2012nu ; Gupta:2011wh . The global structure of the QCD phase diagram, spanning temperature ($T$), baryon chemical potential ($\mu_{B}$), and magnetic field ($B$), remains poorly understood Andersen:2014xxa . At low $\mu_{B}$ and $B=0$, QCD matter undergoes a smooth crossover, while at high $\mu_{B}$, a first-order phase transition occurs. Locating the critical endpoint (CEP) separating these regimes is a primary goal of heavy-ion collision experiments Aoki:2006we ; HotQCD:2018pds .
Magnetic fields, generated in non-central heavy-ion collisions and present in astrophysical objects like magnetars, significantly influence the QCD equation of state. Strong magnetic fields modify the critical temperature, induce novel phases, and play a key role in shaping the QCD phase structure Andersen:2014xxa , making $B$ an essential factor in searching CEP experimentally STAR:2020tga STAR:2021fge STAR:2022hbp . Exploring the magnetic QCD phase structure remains a significant challenge.
With substantial progress, current non-perturbative approaches to studying the QCD phase diagram still face significant challenges. As a first-principle approach, Lattice QCD encounters the well-known sign problem at finite baryon density, making numerical simulations at high $\mu_{B}$ computationally prohibitive Philipsen:2012nu . Moreover, simulating high magnetic field configurations demands substantial computational resources, leading to limited exploration of the QCD phase structure in these regions Andersen:2014xxa . These limitations significantly hinder our understanding of the interplay among $T$, $\mu_{B}$, and $B$, particularly in the regions of the phase diagram where experimental observations are most challenging.
As a competitive approach, holography offers a computationally efficient framework to study non-perturbative QCD systems by mapping them to classical gravity Demircik:2021zll ; Chen:2022goa ; Rougemont:2023gfz . The holographic framework has particularly effectively captured QCD dynamics at finite temperatures and densities Gubser:2008yx ; DeWolfe:2010he ; DeWolfe:2011ts . Several refined models, only capturing corner of the QCD phase diagram, still fall short of quantitatively describing the full set of lattice QCD data, even for QCD equations of state Cai:2012xh ; He:2013qq ; Alho:2013hsa ; Critelli:2017oub ; Knaute:2017opk ; Gursoy:2017wzz ; Yang:2020hun ; Grefa:2021qvt . To construct a holographic model consistent with QCD data, one must systematically explore various actions and scan the parameter space—a task akin to finding a needle in a haystack. Manual parameter tuning to fit lattice QCD data, known as the inverse problem, is computationally inefficient and often fails to capture the system’s complexity, particularly in the multi-dimensional parameter space of $T$, $\mu_{B}$, and $B$ Finazzo:2016mhm ; Rougemont:2023gfz ; Gursoy:2017wzz .
Neural ordinary differential equations (neural ODEs) excel in modeling complex, continuous dynamic systems by optimizing over infinite-dimensional parameter spaces with adaptive precision. We leverage neural ODEs in holography to address these challenges, offering a data-driven mechanism that surpasses traditional machine learning methods relying on empirical formulas for finite-dimensional parameter fitting Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2020jug ; Hashimoto:2022eij ; Chen:2024ckb . This innovative approach captures the intricate dependencies of thermodynamic quantities on $T$, $\mu_{B}$, and $B$ with unprecedented accuracy and flexibility footnote .
These advantages allow our approach to a concrete global structure of the phase diagram and uncover unprecedented novel features of the QCD phase structure, such as multiple CEPs and their associated critical behaviors, which were previously never identified. In particular, for sufficiently large $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, the $T$-$\mu_{B}$ plane exhibits two CEPs, a feature exceeding prior expectations. Furthermore, we analyze the critical behavior near the CEPs, finding that the critical exponents satisfy scaling relations but vary based on the CEP’s location. This innovative combination of machine learning and holography significantly advances our understanding of the QCD phase structure and provides new avenues for better agreement with experimental and lattice QCD results.
Holographic model–To capture essential QCD dynamics at finite magnetic field, temperature, and baryon chemical potential, we employ a holographic framework based on the five-dimensional gravitational theory:
$S={1\over 2\kappa_{N}^{2}}\int d^{5}x\sqrt{-g}\Big{[}\mathcal{R}-{1\over 2}\nabla_{\mu}\phi\nabla^{\mu}\phi$
$-{Z(\phi)\over 4}F_{\mu\nu}F^{\mu\nu}-{\hat{Z}(\phi)\over 4}\hat{F}_{\mu\nu}\hat{F}^{\mu\nu}-V(\phi)\Big{]}\,,$ (1)
where $\kappa_{N}^{2}$ is the effective Newton constant. The metric $g_{\mu\nu}$ characterizes spacetime geometry, and the real scalar field $\phi$ accounts for conformal symmetry breaking. The Maxwell field $A_{\mu}$ with $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ introduces a finite baryon number density, while the magnetic field $B$ is described by another Maxwell field $\hat{A}_{\mu}$ with $\hat{F}_{\mu\nu}=\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$. The functions $Z(\phi)$, $\hat{Z}(\phi)$, and $V(\phi)$ that encode the non-perturbative features of our system are calibrated against lattice QCD data.
The magnetic field $B$ breaks Lorentz invariance along the $z$-axis, leading to anisotropic pressure. Solving the bulk equations of motion (EoMs) allows us to extract thermodynamic quantities, such as the free energy density $\Omega$, longitudinal pressure $P_{z}$, entropy density $s$, baryon density $n_{B}$, and magnetization $M$. For further details, see Supplementary Material Supplementary-Material . This holographic approach provides a first-principles description of QCD dynamics under extreme conditions.
Figure 1: Schematic of the neural ODE approach for solving the inverse problem in QCD phase diagram analysis. Initial conditions from the black hole (BH) horizon are connected to QCD boundary data via EoMs. The neural network initialized as a trial function for the coupling $\hat{Z}(\phi)$ is optimized via backpropagation to minimize the loss function L, representing the difference from lattice QCD data. Parameters $\xi$ are optimized using gradient descent. Technical details are provided in the Supplementary Material Supplementary-Material .
Figure 2: Thermodynamic Quantities from Holographic QCD Model vs. Lattice Data. Temperature dependence of (a) magnetic susceptibility $\chi_{B}$, (b) magnetization $M$, (c) entropy density $s/T^{3}$, and (d) longitudinal pressure $\Delta p_{z}=p_{z}|_{B}-p_{z}|_{B=0}$ across magnetic fields. Shaded regions show lattice QCD estimates Bali:2014kia ; solid lines indicate model predictions. Here, $e=1$, giving $B=1\,\text{GeV}^{2}=1.602\times 10^{19}$ Gauss.
Neural ODEs—In the absence of a first-principle method to determine the coupling functions in our bottom-up model, we constrain these functionals using available lattice QCD data—a challenge known as the inverse problem. As lattice QCD data at finite $B$ and $\mu_{B}$ become more abundant, manually tuning control parameters becomes impractical. Instead of traditional neural networks that approximate solutions in a fixed finite-dimensional parameter space Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2020jug ; Hashimoto:2022eij ; Chen:2024ckb , our neural ODE approach systematically explores the infinite-dimensional parameter space of the coupling functions, achieving higher precision and accuracy.
In practice, we impose boundary conditions at the ultraviolet (UV) boundary and the black hole event horizon. By solving the EoMs numerically, the neural ODE generates a trial equation of state that is iteratively optimized through backpropagation to reduce deviations from lattice QCD data. This process efficiently converges to an optimal magnetic coupling, providing a good agreement with lattice data and refining model predictions across unexplored magnetic fields and chemical potentials (see Fig. 1 for illustration). A detailed algorithmic breakdown of this methodology is included in the Supplementary Material Supplementary-Material . We construct a (2+1)-flavor holographic QCD (hQCD) model using this neural ODE approach to achieve a precise fit with lattice QCD results Bali:2014kia . A similar neural ODE-based method is applied to determine $Z(\phi)$ and $V(\phi)$ by matching to data at $B=0$ HotQCD:2014kol ; Borsanyi:2021sxv . The resulting functional forms are benchmarked rigorously and agree with recent lattice simulations and experimental data Cai:2022omk ; Li:2023mpv .
Fig. 2 presents our holographic predictions for four independent thermodynamic quantities: magnetic susceptibility $\chi^{B}$, magnetization $M$, entropy density $s$, and longitudinal pressure $p_{z}$, compared to lattice QCD data Bali:2014kia . We find good agreement across the available magnetic fields, supporting our holographic model. This is the first holographic model to achieve good agreement with lattice data for magnetic fields up to $B=0.6\text{ GeV}^{2}$. Further analysis is provided in the Supplementary Material, showing consistency with lattice QCD results Bali:2014kia .
QCD phase diagram—With the model fully established, we construct the QCD phase diagram at finite $B$, $T$, and $\mu_{B}$ by computing the free energy density $\Omega$. The full phase diagram is depicted in Fig. 3. The light blue area denotes the first-order phase transition surface, dividing the quadrant into two parts: the high-temperature region corresponds to the quark-gluon plasma, while the low-temperature region corresponds to the hadron gas phase. The deep blue line in the diagram marks the location of CEP for various magnetic fields, where the first-order phase transition terminates and transitions into a smooth crossover at small chemical potentials. At $B=0$, the phase diagram was presented in Fig. 3 of Cai:2022omk , where the first-order transition line terminates at $(T_{C}=105\,\text{MeV},\mu_{C}=555\,\text{MeV})$. Similarly, the phase structure in the $T$-$B$ plane at $\mu_{B}=0$ reveals a line of first-order transitions ending at the CEP with $(T_{C}=89.6\,\text{MeV},B=1.6\,\text{GeV}^{2})$, consistent with lattice QCD predictions Cuteri:2023evl .
Fig. 3 highlights the following three key observations longversion .
  1. 1.
    As the magnetic field $B$ increases up to $B=1.764\,\text{GeV}^{2}$, the critical chemical potential $\mu_{C}$ at the CEP decreases, indicating that stronger magnetic fields shift the CEP to lower chemical potentials. This shows a significant effect of magnetic fields on the CEP location, affecting conditions in experimental settings like heavy-ion collisions.
  2. 2.
    The critical temperature $T_{C}$ at the CEP initially decreases with increasing $B$, reaching a minimum before increasing again. This turning point occurs around $T=80\text{MeV}$, $B=1.6\sim 1.7\text{GeV}^{2}$, and $\mu_{B}=0.2\sim 0.28\text{GeV}$. This behavior suggests complicated effects in the presence of a background magnetic field. It could related to the inverse magnetic catalysis and magnetic catalysis reported in the literature.
  3. 3.
    Multiple CEPs emerge at sufficiently strong magnetic fields in the $T$-$\mu_{B}$ phase diagram, as shown in Fig. 4. A first-order phase transition occurs for $0<\mu_{B}<\mu_{C1}$ and $\mu_{B}>\mu_{C2}$, while a crossover exists when $\mu_{C1}<\mu_{B}<\mu_{C2}$. As the magnetic field $B$ increases, the two CEPs, initially distinct, converge into a single point. Specifically, Fig. 4 illustrates two CEPs identified at $(T_{C},\mu_{C})=(87.3\,\mathrm{MeV},115.9\,\mathrm{MeV})$ and $(T_{C},\mu_{C})=(78.9\,\mathrm{MeV},244.0\,\mathrm{MeV})$. This result holds significant implications for heavy-ion collision experiments, providing precise experimental markers for testing at FAIR, JPARC-HI, and NICA Fukushima:2020yzx .
These key observations reveal a rich phase structure in a strong magnetic field and warrant further experimental verification.
Figure 3: QCD phase diagram at finite magnetic field $B$. Phase structure in temperature $T$, baryon chemical potential $\mu_{B}$, and magnetic field $B$ from our holographic model. The light blue surface denotes the first-order transition boundary, separating the hadronic phase from the quark-gluon plasma. The dark blue line traces the CEP trajectory, marking where the first-order transition ends in a crossover.
Figure 4: QCD phase diagram in the $T$-$\mu_{B}$ plane at $B=1.618\,\mathrm{GeV}^{2}$. The purple line marks the first-order phase transition, ending at the first critical endpoint (CEP) at $\mu_{C1}=115.9$ MeV, where the transition becomes a crossover. A second CEP at $\mu_{C2}=224.0$ MeV suggests a more complex phase structure. These findings imply that strong magnetic fields induce rich QCD phase behavior, with significant implications for heavy-ion collisions.
Critical exponents—Beyond mapping the phase diagram, we examine critical behavior near the CEPs via critical exponents. These exponents describe how thermodynamic quantities, such as susceptibility and specific heat, diverge near critical points, typically following power-law scaling. These exponents help identify the CEP’s universality class and provide insights into QCD transitions under extreme conditions.
Four critical exponents can be directly extracted from the phase diagram of Fig. 3.
  • Critical exponent $\alpha$: The exponent $\alpha$ quantifies the power-law behavior of specific heat near a CEP along the axis defined as approaching the CEP along the tangent of the first-order line:
    $C_{n}=T\left({\partial s\over\partial T}\right)_{n_{B},B}\sim|T-T_{\text{CEP}}|^{-\alpha}.$
  • Critical exponent $\beta$: It characterizes the discontinuity of entropy density $s$ across the first-order line:
    $\Delta s=s_{>}-s_{<}\sim(T_{\text{CEP}}-T)^{\beta},$
    where $s_{>}$ and $s_{<}$ represent the entropy densities in the high- and low-temperature phases, respectively.
  • Critical exponent $\gamma$: It represents the power-law behavior of baryon susceptibility with the temperature near the CEP along the first-order axis:
    $\chi_{2}^{B}={1\over T^{2}}\left({\partial n_{B}\over\partial\mu_{B}}\right)_{T,B}\sim|T-T_{\text{CEP}}|^{-\gamma}.$
  • Critical exponent $\delta$ : The definition of $\delta$ relies on the power-law relationship between entropy and chemical potential with $T=T_{\text{CEP}}$ at the critical isotherm:
    $s-s_{\text{CEP}}\sim|\mu_{B}-\mu_{B\text{CEP}}|^{1/\delta},$
    where $s_{\text{CEP}}$ is the entropy density at the CEP.
Table 1 presents the critical exponents for the CEP at three different magnetic field values, denoted by hQCD (I, II, III) longversion . The critical exponents satisfy the scaling relations $\alpha+2\beta+\gamma=2$ and $\alpha+\beta(1+\delta)=2$, ensuring the self-consistency of our results. Although close to mean-field values, these exponents show significant deviations depending on the CEP location, particularly as $B$ increases. Such deviations highlight the features of our holographic QCD model, which cannot be attributed to large-$N$ effects typical in conventional holographic duality, where mean-field behavior is expected. Similar deviation happened in the holographic 2-flavor model Zhao:2023gur , the critical exponents match those of the quantum 3D Ising model, further emphasizing the distinct nature of our approach compared to traditional large-$N$ QCD models.
$\alpha$ $\beta$ $\gamma$ $\delta$
Experiment 0.110-0.116 0.316-0.327 1.23-1.25 4.6-4.9
3D Ising 0.110(5) 0.325$\pm$0.0015 1.2405$\pm$0.0015 4.82(4)
Mean field 0 1/2 1 3
DGR model 0 0.482 0.942 3.035
hQCD(I) 0.002296 0.485518 0.9558187 3.00993
hQCD(II) 0.001694 0.50373 0.91803 2.9455
hQCD(III) 0.00917 0.3944 0.98696 3.9878
Table 1: Critical exponents from experiments in non-QCD fluids, the full quantum 3D Ising model, mean-field (van der Waals) theory, the DGR model DeWolfe:2010he , and our (2+1)-flavor hQCD model. The hQCD (I, II, III) correspond to the critical exponents for $\mu_{B}=554.66\ \text{MeV},B=0$ (hQCD I), $\mu_{B}=501.4\ \text{MeV},B=0.3\ \text{GeV}^{2}$ (hQCD II), and $\mu_{B}=0,B=1.6\ \text{GeV}^{2}$ (hQCD III), respectively.
Conclusion– We have developed a novel neural ODE framework that solves the inverse problem of constructing a holographic QCD action from Lattice QCD data. This framework results in the first holographic model capable of capturing key thermodynamic behaviors of hot and dense QCD at finite magnetic fields. Notably, the model reveals a rich phase structure in a strong magnetic field, including non-monotonic CEP temperature behavior (Fig. 3) and multiple CEPs in the $T$-$\mu_{B}$ plane (Fig. 4), providing specific experimental markers for validation at future facilities like FAIR, JPARC-HI, and NICA Fukushima:2020yzx . Specifically for $B=1.618\text{GeV}^{2}$, we identified the first CEP at ($T_{C}=87.3,\mathrm{MeV},\mu_{C}=115.9,\mathrm{MeV}$), and a second CEP at ($T_{C}=78.9,\mathrm{MeV},\mu_{C}=244.0,\mathrm{MeV}$). These results suggest a richer phase structure in QCD than previously anticipated. Furthermore, we have determined critical exponents that depend on the location of CEPs, offering valuable insights for experimental studies in regions accessible to RHIC and LHC. Experimental observables, such as baryon number or magnetization fluctuations, could directly test these predictions in current experiments, e.g. RHIC STAR:2020tga , the STAR fixed target program (FXT), and future experiments Fukushima:2020yzx .
Future research should extend the model to bridge gaps between theoretical predictions and experimental findings across energy scales. This includes incorporating isospin asymmetry, which is relevant for neutron stars, and rotational effects, which are critical for understanding dynamics in rapidly spinning neutron stars and heavy-ion collisions. Extending the model to non-equilibrium scenarios could also provide insights into the real-time dynamics of phase transitions, capturing rapid changes in temperature, density, and magnetic field during heavy-ion collisions. Applying our findings to neutron stars and early universe conditions, where understanding the equation of state for strongly interacting QCD matter under varying magnetic fields is crucial, presents another promising direction for future research.
Acknowledgements–We thank Heng-Tong Ding, Matti Järvinen, Elias Kiritsis, Zhibin Li, Xiaofeng Luo, Yi-Bo Yang, Danning Li, Javier Subils, Ling-Xiao Wang, and Yuan-Xu Wang for stimulating discussions. This work is supported in part by the National Key Research and Development Program of China Grants No.2020YFC2201501 and No.2020YFC2201502, in part by the National Natural Science Foundation of China Grants No.12075101, No.12047569, No.12122513, No.12075298, No.11991052, No.12235016 and No.12047503. S.H. would also like to express appreciation for the financial support from the Max Planck Partner Group.